OpenStudy (anonymous):

(3+square root 6)(3-square root 6)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[(3+\sqrt{6}) \times (3-\sqrt{6})\]

OpenStudy (anonymous):

can you simplify from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

does it want you to simplify or solve?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

Ok then, let me think.

OpenStudy (anonymous):

use FOIL - first, outside, inside, last. (for binomials) \[(6 -3\sqrt{6} + 3\sqrt{6} + -\sqrt{12})\]

OpenStudy (anonymous):

Combine like terms \[(6 -\sqrt{6} -\sqrt{12})\]

OpenStudy (anonymous):

And i can't remember how to simplify from there sorry >.<

OpenStudy (anonymous):

@SolomonZelman can you help?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf (a+b)(a-b)=a^{2}-b^2}\) your "a" is 3, and your "b" is √6 .

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

now, @iminlove_123 please show me the steps.

OpenStudy (anonymous):

im cinfused

OpenStudy (anonymous):

Substitute 3 for "a" and sqrt{6} for "b"

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf (3-\sqrt{6}) (3+\sqrt{6})=3^{2}-(\sqrt{6}) ^{2}=9-6=3 }\)

OpenStudy (anonymous):

oh ok i get it

OpenStudy (anonymous):

you would get \[(3 + \sqrt{6}) \times (3 - \sqrt{6}) = 3^{2} - \sqrt{6}^{2}\]

OpenStudy (anonymous):

Dang message me back